OpenStudy (anonymous):

how do u solve: 4m-14/7-2m what are the steps

OpenStudy (anonymous):

use the same Technic as in precious example: =2*(2m-7)/-(2m-7)=-2

OpenStudy (anonymous):

You have to recognize that 4m - 14 = -2 * (7 - 2m). Then you have -2(7 - 2m)/(7 - 2m) = -2